# When Large Language Models Meet Speech: A Survey on Integration Approaches

### v1

@@ A Survey on Integration Approaches
Zhengdong Yang  Shuichiro Shimizu  Yahan Yu  Chenhui Chu Kyoto University {zd-yang, sshimizu, yahan}@nlp.i.kyoto-u.ac.jp  chu@i.kyoto-u.ac.jp
@@ 1 Introduction
Figure 1: Overview of three fundamental approaches for speech-LLM integration.
@@
SLAM-ASR Ma et al. (2024), Denisov and Vu (2024) , MaLa-ASR Yang et al. (2024b),
@@ 3.2 LLM Rescoring
As shown in Figure 3 (a), this process involves generating a list of $n$-best hypotheses =${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{H}}=\{Y_{1},Y_{2},...,Y_{n}\}$ from the speech input $X$ through the initial ASR decoding, which is then re-evaluated using an LLM to find the hypothesis with higher linguistic coherence to improve accuracy. The rescoring process can be expressed as
$Y^{*}$ $=\text{argmax}_{Y_{i}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{H}}}\ [(1\!-\!\lambda)\log p_{\text{AM}}(Y_{i}|X)$
@@ 4 Latent-representation-based Integration
In this integration approach, a speech encoder is used to process the speech input and generate latent representations that are directly fed into the LLM, bypassing the embedding layer. The speech encoder is usually a Transformer-based model, which can be pre-trained on large-scale speech data or trained from scratch for the speech-LLM integration.
@@ 4.2.3 Q-Former
Many studies Yu et al. (2024); Tang et al. (2024); Pan et al. (2023) adopt Q-Former for the modality adaptation. Q-Former is a Transformer-based module converting variable-length input sequences into fixed-length output query representations. It is initially proposed for visiontext modality alignment Li et al. (2023a).
@@ 5.1.3 Integration of Semantic and Acoustic tokens
Task Dataset Metrics Model Integration Method Performance
ASR Librispeech dev-clean | dev-other | test-clean | test-other WER $\downarrow$ Whisper-large-v2 Radford et al. (2023) Non-LLM — | — | 2.7 | 5.2
@@
Fleurs zh | en WER $\downarrow$ Whisper-large-v2 Radford et al. (2023) Non-LLM 4.2 | 14.7
@@
AISHELL-2 Mic | iOS | Android WER $\downarrow$ Seed-ASR Bai et al. (2024b) Latent-representation-based $\rightarrow$ Convolutional Downsampling 2.2 | 2.2 | 2.2
@@
VoxPopoli All WER $\downarrow$ SpeechVerse Das et al. (2024) Latent-representation-based $\rightarrow$ Convolutional Downsampling 6.5
@@
S2TT CoVoST2 en-de | de-en | en-zh | zh-en | es-en | fr-en | it-en | ja-en BLEU $\uparrow$ Whisper-large-v2 Radford et al. (2023) Non-LLM — | 36.3 | — | 18.0 | 40.1 | 36.4 | 30.9 | 26.1
@@
S2ST CVSS S2ST de-en | zh-en | es-en | fr-en | it-en | ja-en ASR-BLEU $\uparrow$ Translatotron 2 + pretraining
@@
TTS AISHELL-1 CER $\downarrow$ | SECS $\uparrow$ | MOSNet $\uparrow$ VALL-E Phone Wang et al. (2023b) Non-LLM 4.75 | 0.91 | 3.22
@@
LibriTTS WER $\downarrow$ | SECS $\uparrow$ | MOSNet $\uparrow$ VALL-E Phone Wang et al. (2023b) Non-LLM 4.30 | 0.92 | 3.28
@@
Topic-StoryCloze TSC $\uparrow$ Spirit-LM Nguyen et al. (2025) Audio-token-based $\rightarrow$ Semantic Token 82.9
@@
StoryCloze SSC $\uparrow$ Spirit-LM Nguyen et al. (2025) Audio-token-based $\rightarrow$ Semantic Token 61.0
@@ 6.1 Advantages and Disadvantages
  • Degree of integration: Latent-representation-based $>$ Audio-token-based $>$ Text-based.
  • Interpretability: Text-based $>$ Audio-token-based $>$ Latent-representation-based.
@@ 6.2 Quantitative Comparison
We provide a comprehensive performance comparison in Table 1, comparing studies with different integration approaches across multiple tasks: ASR, S2TT, S2ST, TTS. The table highlights key metrics such as word error rate (WER), BLEU score, and others, alongside each model’s integration method and backbone components.
@@ 7.2 Latent-representation-based Integration
In contrast to text-based integration, latent-representation-based integration employs an intermediate representation that is closer to the source speech but more distant from the LLM’s natural input space. Consequently, the primary challenge lies in aligning these acoustic representations with the textual embedding space of the LLM. Section 4.2 introduces multiple modality adaptation mechanisms to address this and enable LLMs to handle speech more effectively.
While the primary focus of most research introduced in this paper has been on adapting LLMs to better handle speech data, there is comparatively less attention on how these adaptations affect performance on text-based tasks. The alignment of text and speech data, along with their latent representations, remains underexplored, with only a few studies focusing on this issue Wang et al. (2023a); Lu et al. (2024).
@@ 7.4 Fair Comparison Across Integration Approaches
Beyond these integration-approach-specific challenges, there is a notable gap in comparing the different integration approaches under a unified setting. Most existing works focus on one approach, making it difficult to assess their relative merits consistently. A fair comparison among these integration methods could clarify how different factors affect performance. Developing standardized benchmarks, protocols, and reporting practices for speech-LLM research would help future work isolate the core differences between these approaches.

### v2

@@ A Survey on Integration Approaches
Zhengdong Yang, Shuichiro Shimizu, Yahan Yu, Chenhui Chu Kyoto University {zd-yang, sshimizu, yahan}@nlp.ist.i.kyoto-u.ac.jpchu@i.kyoto-u.ac.jp
@@ 1 Introduction
Figure 1: Overview of three fundamental approaches for speech–LLM integration.
@@
SLAM-ASR Ma et al. (2024), Denisov and Vu (2024a) , MaLa-ASR Yang et al. (2024b),
@@ 3.2 LLM Rescoring
As shown in Figure 3 (a), this process involves generating a list of $n$-best hypotheses $\mathcal{H}=\{Y_{1},Y_{2},...,Y_{n}\}$ from the speech input $X$ through the initial ASR decoding, which is then re-evaluated using an LLM to find the hypothesis with higher linguistic coherence to improve accuracy. The rescoring process can be expressed as
$Y^{*}$ $=\text{argmax}_{Y_{i}\in\mathcal{H}}\ [(1\!-\!\lambda)\log p_{\text{AM}}(Y_{i}|X)$
@@ 4 Latent-representation-based Integration
In this integration approach, a speech encoder is used to process the speech input and generate latent representations that are directly fed into the LLM, bypassing the embedding layer. The speech encoder is usually a Transformer-based model, which can be pre-trained on large-scale speech data or trained from scratch for the speech–LLM integration.
@@ 4.2.3 Q-Former
Many studies Yu et al. (2024); Tang et al. (2024); Pan et al. (2023) adopt Q-Former for the modality adaptation. Q-Former is a Transformer-based module converting variable-length input sequences into fixed-length output query representations. It is initially proposed for vision–text modality alignment Li et al. (2023a).
@@ 5.1.3 Integration of Semantic and Acoustic tokens
Task
Dataset
Metrics Model Integration Method Performance
ASR
Librispeech
dev-clean | dev-other |
test-clean | test-other
WER $\downarrow$ Whisper-large-v2 Radford et al. (2023) Non-LLM — | — | 2.7 | 5.2
@@
Fleurs
zh | en
WER $\downarrow$ Whisper-large-v2 Radford et al. (2023) Non-LLM 4.2 | 14.7
@@
AISHELL-2
Mic | iOS | Android
WER $\downarrow$ Seed-ASR Bai et al. (2024b) Latent-representation-based $\rightarrow$ Convolutional Downsampling 2.2 | 2.2 | 2.2
@@
VoxPopoli
All
WER $\downarrow$ SpeechVerse Das et al. (2024) Latent-representation-based $\rightarrow$ Convolutional Downsampling 6.5
@@
S2TT
CoVoST2
en-de | de-en |
en-zh | zh-en |
es-en | fr-en |
it-en | ja-en
BLEU $\uparrow$ Whisper-large-v2 Radford et al. (2023) Non-LLM — | 36.3 | — | 18.0 | 40.1 | 36.4 | 30.9 | 26.1
@@
S2ST
CVSS S2ST
de-en | zh-en |
es-en | fr-en |
it-en | ja-en
ASR-BLEU $\uparrow$ Translatotron 2 + pretraining
@@
TTS
AISHELL-1
CER $\downarrow$ | SECS $\uparrow$ |
MOSNet $\uparrow$
VALL-E Phone Wang et al. (2023b) Non-LLM 4.75 | 0.91 | 3.22
@@
LibriTTS
WER $\downarrow$ | SECS $\uparrow$ |
MOSNet $\uparrow$
VALL-E Phone Wang et al. (2023b) Non-LLM 4.30 | 0.92 | 3.28
@@
Topic-StoryCloze
TSC $\uparrow$
Spirit-LM Nguyen et al. (2025) Audio-token-based $\rightarrow$ Semantic Token 82.9
@@
StoryCloze
SSC $\uparrow$
Spirit-LM Nguyen et al. (2025) Audio-token-based $\rightarrow$ Semantic Token 61.0
@@ 6.1 Advantages and Disadvantages
  • Degree of integration: Latent-representation-based >> Audio-token-based >> Text-based.
  • Interpretability: Text-based >> Audio-token-based >> Latent-representation-based.
@@ 6.2 Quantitative Comparison
We provide a comprehensive performance comparison in Table 1, comparing studies with different integration approaches across multiple tasks: ASR, S2TT, S2ST, and TTS. The table highlights key metrics such as word error rate (WER) and BLEU, alongside each model’s integration method and backbone components.
@@
6.3 Computational Trade-offs
As different methods have different computational requirements, it is difficult to have direct comparisons through a literature review. For numerical comparisons, comparing different approaches with a standard setting is required, which is one future direction. That said, we provide an overview of the computational requirements per category below:
  • For text-based integration, the computational cost is equivalent to that of running separate models for ASR, LLM, and TTS. Rescoring based on $N$-best list increases the inference compute by a constant factor. For generative error correction, studies have explored full fine-tuning, parameter-efficient fine-tuning, and mixture-of-experts of LLMs. Parameter-efficient fine-tuning can provide better performance compared to full fine-tuning Chen et al. (2023a). Mixture-of-experts requires high memory and computation for training, but shows good performance in handling multiple tasks with light inference.
  • For latent-representation-based integration, speech encoders are connected to LLMs through adapters. These adapters are lightweight, and studies have explored full fine-tuning and parameter-efficient fine-tuning, including freezing a part of the models. Pham et al. (2024) showed that applying LoRA to the LLM significantly enhances the performance, and for the encoder module, full fine-tuning yields the best performance, followed by partial fine-tuning, while partial fine-tuning is more cost-effective.
  • For audio-token-based integration, it involves audio token encoding/decoding and language modeling with audio tokens. For semantic tokens, feature extraction from S3Ms and training/assignment steps of k-means is required. They are lightweight compared to language modeling (the largest S3Ms have about 2B parameters), but efficient implementation is required for large-scale data. If the output is speech, a vocoder to convert semantic tokens to waveforms is required, which is also relatively lightweight compared to language modeling. For acoustic tokens, encoding and decoding are done using the same model, and the inference is lightweight. Computational cost for language modeling is similar to that of text-based ones.
Overall, we can argue that the computational cost is bottlenecked by the size of the LLM that is used. Text-based integration and audio-token-based integration would require roughly the same computational bottleneck by LLMs, while latent-representation-based integration requires handling both the speech encoder and LLMs at the same time.
@@ 7.2 Latent-representation-based Integration
In contrast to text-based integration, latent-representation-based integration employs an intermediate representation closer to the source speech but more distant from the LLM’s native input space. Consequently, the primary challenge lies in aligning these acoustic representations with the textual embedding space of the LLM. Section 4.2 introduces multiple modality adaptation mechanisms to address this and enable LLMs to handle speech more effectively.
While the primary focus of most research introduced in this paper has been on adapting LLMs to handle speech data better, there is comparatively less attention on how these adaptations affect performance on text-based tasks. The alignment of text and speech data, along with their latent representations, remains underexplored, with only a few studies focusing on this issue Wang et al. (2023a); Lu et al. (2024).
@@ 7.4 Fair Comparison Across Integration Approaches
Beyond these integration-approach-specific challenges, there is a notable gap in comparing the different integration approaches under a unified setting. Most existing works focus on one approach, making it difficult to assess their relative merits consistently. A fair comparison among these integration methods could clarify how different factors affect performance. Developing standardized benchmarks, protocols, and reporting practices for speech–LLM research would help future work isolate the core differences between these approaches.
7.5 Multilingualism
The challenge of multilingualism in speech–LLM integration is also significant. Most existing speech–LLM models are designed to handle only English, which limits their range of applicability. Although there are studies focusing on multilingual capabilities Hu et al. (2024c, 2023); Huang et al. (2024b), the majority still rely on English-centric LLMs. Only a small fraction of research efforts Denisov and Vu (2024b) explicitly consider the multilingualism of pre-trained LLMs. However, the effectiveness of a speech–LLM model is often constrained by the quality of the text-based backbone LLM. This dependency implies that any limitations in the backbone LLMs, especially in terms of handling multiple languages, can directly impact the performance of speech LLMs.
7.6 Real-time Processing
Real-time processing presents another substantial challenge in the field of speech–LLM integration. Compared to text-based applications, speech-based applications such as ST more often require real-time processing to be practical in everyday use. However, the large size of LLMs typically leads to increased latency, which can neutralize some of the benefits gained from improved accuracy. While most speech LLMs are capable of handling tasks offline, there is a need for more innovative methods that can facilitate online processing. Deeper integration with latent representations or audio tokens tends to result in lower latency compared to text-based approaches. Recent models that focus on real-time applications, such as Moshi Défossez et al. (2024), show promising starts in addressing this challenge. Developing such methods would significantly enhance the utility of LLMs in real-time speech applications, making them more viable for real-world usage.
@@ 8 Conclusion
Acknowledgment
This work was supported by the SPRING program of Kyoto University and JSPS KAKENHI Grant Number JP23K28144 and JP23KJ1356.